OpenStudy (anonymous):

the sum of the height and diameter of a cylinder is 24 cm. what is the maximal volume of the cylinder?

OpenStudy (anonymous):

\[22/7\times23\times1\]

OpenStudy (anonymous):

how did u get the answer please ?

OpenStudy (anonymous):

You can find the maximum volume by differentiation. Find the critical point, V'(r)=0 well, \[2\pi r+h=24cm\] therefore, \[h=24-2 \pi r\] and the volume of the cylinder is\[v(r)=\pi r^2h=\pi r^2(24-2 \pi r)\] you differentiate this, \[v'(r)=48 \pi r-6 \pi^2 r^2=0\] Therefore, \[\pi r=8\](or 0 but it is unreasonable cause neither r and pi is not 0) So,\[2 \pi r=16cm, h=8cm, r=2.55cm\] The answer is v=3.14*(2.55)^2*8=163.34cm^3 But I don't know this is right or wrong. Sorry :)

OpenStudy (anonymous):

thank you